Grade 12th passMechanics

Two masses of 2 kg and 4 kg are held with a compressed spring between them. If the masses are released, the spring will push them away from each other. If the smaller mass moves off with a velocity of 6m/s, what is the stored energy in the spring when it is compressed? Asume there is no external force.

Mexiprist
8 Years agoGrade 12th pass

2 Answers

Sujan Chegu
8 Years ago
By conservation of momentum, 2*6=4*vV=3.Net kinetic energy finally is 36+18=54 JSo intial spring energy is 54J
Sujan Chegu
8 Years ago
By conservation of momentum, 2*6=4*vV=3.Net kinetic energy finally is 36+18=54 JSo intial spring energy is 54J.